OpenStudy (anonymous):

62 TIMES WHAT WOULD EQUAL 367 I PROMISE WILL GIVE MEDAL 2 FIRST RESPONSE

OpenStudy (benlindquist):

hold on

OpenStudy (benlindquist):

16

OpenStudy (christos):

62*x=367

OpenStudy (christos):

x = 367/62

OpenStudy (anonymous):

62x = 367 divide both sides by 62 x = 367/62 x = 5.91935

OpenStudy (benlindquist):

idk

OpenStudy (anonymous):

ALREADY GOT MY BEST RESPONSE THANKS FOR YOUR HELP BEHIND QUIST

OpenStudy (christos):

so 62 times 367/62 would equal 367

OpenStudy (anonymous):

@kennyboy00 you do know he's wrong tho right... 62 * 16 is 992

OpenStudy (mrnood):

@kennyboy00 @benlindquist is WRONG

OpenStudy (benlindquist):

i know do be like he's an idiot

OpenStudy (anonymous):

MR NOOD IF HES WRONG CAN u HELP MEEE